OpenStudy (anonymous):

what do I do after this step? there are 24 coins, some nickels and some quarters, which add up to $3.40. ok, so ik to do n+q=24 and .05n+.25q=3.4, but idk what to do next

OpenStudy (anonymous):

solve the 2 eqns n+q=24------eqn(1) .05n+.25q=3.4 -----eqn(2) multiply above eqn with 100 5n+25q=340 -------eqn(III) we need to solve eqn (I) and(III) multiply eqn (I) with 5 n+q=24 5n+5q=120----eqn(IV)

OpenStudy (anonymous):

solve eqn (III) and (IV) subtract eqn(IV) from eqn(III) 5n+25q=340 -(5n+5q=120) -------------- 20q=220 q=11 put value of q in any of given eqn n+q=24 n+11=24 n=13 so n=13 and q=11